Math 213a: Complex analysis Handout #0 (15 September 2003): About Math 213a

What? Math 213a is a graduate-level introduction to the classical theory of analytic functions of one complex variable from a modern viewpoint. The main topics are as listed in the Course Catalog, and in particular include all the complex analysis that graduate students in mathematics will need for their Qualifying Examination.

Who? As is true of all 200-level courses, Math 213a is primarily intended for graduate students; undergraduates generally learn complex analysis in Math 113. However, advanced undergraduates may enroll in 213, and each year several math majors do this. As usual for 200-level courses, I must sign the study card of any undergraduate taking this class for credit.

Textbook The course will loosely follow Ahlfors’ excellent book, of which there will be a few copies on reserve at the Birkhoff (3rd floor math) library and several copies available for perusal or loan at Cabot library (1st floor of the Science Center). You are not required to purchase a copy, but the Coop has some on sale in case you decide to acquire this book.

Grading If you are taking 213a and are not an EXCused graduate student, you will get a grade for the class. Most of your grade (about 2/3) will be based on regular problem sets. There is no final exam; instead, the final 1/3 of your grade will depend mostly on a final project, with class participation used mostly to decide borderline cases. The nature of the final project (expository paper or in-class presentation) will be determined once the class size has stabilized after “Shopping Period”.

Office information, etc. My office is Room 335 of the Science Center (right outside the math library on the 3rd floor), telephone #(49)5-4625; my e-mail address is elkies@math. Office hours: Thursdays 3:00–4:30 (except Colloquium Thursdays, 3:00–4:00), or by appointment. Course URL:

www.math.harvard.edu/~elkies/M213a.03; most of the handouts, prob-

lem sets, etc. for the course will be posted on that page. The CA for Math 213a is Matthew Bainbridge (matt@math). Section place and time will be determined once the class roster has stabilized and we know what everybody’s schedule is.

Note I’ll have to miss a few classes during the course of the term: September 24 and 26, October 6, and maybe once or twice afterwards. These will be announced in advance, and some makeup meetings will be scheduled (probably at the same MWF 12–1 time) during Reading Period.